# On the attribution of weather events to climate change using a fit to extreme value distributions

### v2

@@ Abstract
Increases in extreme weather events are a potentially important consequence of anthropogenic climate change (ACC), yet, are difficult to attribute to ACC because the record length is often similar to, or shorter than, extreme-event return periods. This study is motivated by the “World Weather Attribution Project” (WWA) and, specifically, their approach of fitting extreme value distribution functions to local observations. The approach calculates the dependence of distribution parameters on global mean surface temperature (GMST) and uses this dependence to attribute extreme events to ACC. Applying this method to preindustrial climate simulations with no time-varying greenhouse gas forcing, we still find a strong dependence of distribution parameters on GMST. This dependence results from internal climate variability (e.g., ENSO) affecting both extreme events and GMST. Therefore, dependence on GMST does not necessarily imply an effect of ACC on extremes. For three different cases examined by WWA we find that the parametric relationship between weather and GMST is indistinguishable between simulations that are and are not forced by greenhouse gasses. We further consider for these three cases whether an extreme value, normal, or log-normal distribution better represents the data; if a GMST-dependence of distribution parameters is justified using a likelihood ratio test; and if a meaningful attribution is possible given uncertainties in GMST dependence. In one case we find that an attribution of Australia’s 2020–2021 Bushfires to ACC is difficult due to the effects of Natural variability. In another involving the 2019–2021 drought in Madagascar we find that the small number of available data points precludes a meaningful attribution analysis. Overall, we find that the effects of natural climate variability on GMST and the uncertain relationship between GMST and regional extremes may lead to inaccurate attribution conclusions using the WWA approach.
@@ 1 Introduction
The Intergovernmental Panel on Climate Change Core Writing Team et al. (2023) found that global mean surface temperature (GMST) between 2011–2020 was 1.09C [0.95 C–1.20C] above preindustrial levels. Although climate change is clearly manifested in GMST warming, and there are reasons to expect extreme events to change with mean climate, it is a complex task to attribute individual extreme weather events to anthropogenic climate change. Extreme events—including heatwaves, droughts, and floods—are rare by definition, so the observational record may be too short to detect anthropogenic signals. This issue is especially pervasive in lower-income countries where data are sparse and the historical record is short Otto et al. (2020). The situation is further problematic for extreme precipitation event attribution, which is highly sensitive to the specific timescales and percentiles used to define extreme precipitation events Pendergrass (2018).
@@
The field of attribution of extreme events has progressed significantly over the past decade and encompasses several approaches (see example reviews and cited references within, Stott et al., 2010; Hegerl and Zwiers, 2011; National Academies of Sciences, 2016; Clarke et al., 2022; Philip et al., 2020). The approaches that have been applied include the fraction of attributable risk (FAR, Allen, 2003; Stott et al., 2004), using simulations from climate models, using constraints derived from optimal fingerprinting techniques Christidis et al. (2015), using climate models to validate observed probability density functions (PDFs) Perkins et al. (2007), and the “Analogue-Based Approach” Lorenz (1969); Yiou et al. (2007); Vautard and Yiou (2009); Faranda et al. (2022). Another methodology fits statistical distributions to the historical record (again, either observed or from climate models) in what Stott et al. (2016) refer to as “Empirical-Based Approaches.” This approach assesses how the return period of particular events has changed over the historical record, with the purpose of detecting potential trends in the frequency of extreme events. These statistical approaches have been used to study trends in extreme events such as flooding in Thailand in 2011 Van Oldenborgh et al. (2012). Stott et al. (2016) note that approaches using climate models allow for a better estimate of the magnitude and direction of the human-induced contributions to an extreme event than is possible using observations alone.
@@
Another use of climate models is based on single-model initial condition large ensembles (LE), composed of a climate model run several times with slightly different initial conditions. Averaging over multiple climate model ensemble members can reduce the signal of internal variability, facilitating the identification of forced climate change signals. This averaging generally improves the detection of trends in the magnitude or frequency of an extreme event, which are difficult to quantify in a single simulation or in the observational record where the signal from internal variability may be dominant Yoshida et al. (2017); Yamaguchi et al. (2020). These simulations have been used in the context of tropical cyclone slowdown Yamaguchi et al. (2020), as well as drought events in South and East Africa Lott et al. (2013); Pascale et al. (2020). The detection and attribution literature is extensive, and for a further detailed discussion, we refer the reader to the above reviews and to recent studies Stott et al. (2016); Knutson et al. (2017); Seneviratne et al. (2021).
@@
One particular approach that is the focus of this paper and that has been used in recent years as a rapid response to extreme events attempts to fit the observed data to an extreme value PDF. The fit allows for a dependence of the distribution parameters on explanatory variables, typically the GMST (e.g., Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022; Philip et al., 2020). A GMST dependence is then interpreted as a response of extreme events to anthropogenic global warming. This dependence is then used to attribute extreme events by quantifying the effects of anthropogenic forcing on the probability of their occurrence. Technically this is achieved by writing the location parameter of the extreme event distribution, $\mu$, as a function of $\overline{T}$, the (low-pass filtered, via a 4-year running average) GMST “as a measure of anthropogenic climate change” Harrington et al. (2022), such that $\mu=\mu_{0}+\alpha\overline{T}$. The GMST-dependence, $\alpha$, is estimated via a fit to observations. The fitted PDF is then used to evaluate the probability of extreme events at 1900, $P(\mu_{0}+\alpha\overline{T}(1900))$ relative to those at 2020, $P(\mu_{0}+\alpha\overline{T}(2020))$. This allows for estimating how the probability of an extreme event changed between these times. This approach has been used by the World Weather Attribution Project (WWA, https://www.worldweatherattribution.org/). Whereas our focus is on this observational approach to attribution, it should be noted that WWA also makes use of climate model simulations in interpreting extreme events.
Motivated by these pioneering efforts based on the fit to extreme event PDFs, we re-examine three test cases analyzed by WWA: the Siberian Heatwave analysis of Ciavarella et al. (2021, hereafter CC2021), the analysis of high temperatures associated with the Australian Bushfire by van Oldenborgh et al. (2021, hereafter OK2021), and the Madagascar Drought analysis of Harrington et al. (2022, hereafter HW2022). These cases were chosen as they represent a range of applications, including the attribution of heatwaves at high and low latitudes (Siberia vs. Australia) and of a drought extreme. Each of these studies leverages either a Generalized Extreme Value (GEV) or Generalized Pareto Distribution (GPD) fit to assess potential contributions to extreme events from anthropogenic climate change. WWA studies use both climate model simulations and observational analysis, and we focus on the observational component. In each of the three cases we use the CESM Large Ensemble Kay et al. (2015) to put the observed instrumental record into perspective, and we explore potential caveats concerning the methods examined and suggest how these limitations might be addressed.
An important issue to address before getting into the technicalities is the formulation of the null hypothesis in an attribution study that uses GMST-dependent distribution parameters to deduce the effects of anthropogenic climate change. The effects of anthropogenic climate change on extreme events are parameterized by the term $\alpha$ in WWA studies, $P(\mu_{0}+\alpha\overline{T})$. Although WWA does not provide a detailed discussion of the null hypothesis that they assume in the absence of anthropogenic climate change, the implied null hypothesis is that $\alpha$ is zero (see Section 4.3.1 of Philip et al., 2020).
We find here that in a preindustrial climate model run, the fit of Australian heat extremes leads to values of $\alpha$ that are substantially greater than zero in the absence of anthropogenic climate change because of the effects of natural climate variability (by which we mean internal variability due to modes such as ENSO as opposed to volcanic-driven natural variability). Our findings thus show that the null hypothesis can be rejected even in the absence of anthropogenic climate change, complicating the WWA approach. The issue related to the null hypothesis that our analysis exposes is simple: natural variability modes such as ENSO affect both local temperature extremes and the GMST. In some cases, the effect on local extremes (say the effects of ENSO on heatwaves in Australia) may be large, but the effect on GMST is small. This might lead to the false conclusion that a small GMST signal leads to a comparatively large effect on local extremes.
Our study addresses four questions: (1) Is a GMST dependence of the PDF parameters necessarily an indication of the signal of anthropogenic climate change? (2) Is a GEV or GPD necessary to fit the available observations, or is a more standard distribution (normal/log-normal) sufficient? (3) Is the addition of GMST-dependent distribution parameters, which form the basis for the attribution analysis, justified statistically, and is the error level (the uncertainty range of $\alpha$) in this dependence sufficiently small to allow for attribution? (4) What do Single-Model Initial-Condition Large Ensembles tell us about the amount of data needed to determine the GMST dependence with confidence?
@@
We note that WWA uses multiple attribution tools, including model runs and the analysis of observations Philip et al. (2020). Similarly, WWA often uses multiple climate models and is careful to exclude models based on their performance and return periods relative to the observed one (e.g., Pinto et al., 2023), whereas we use a single model (CESM) to demonstrate some of our points and focus on observational analysis. We focus on one aspect only of this richer approach, involving a fit of observations to extreme value distributions. Additionally, our goal is limited to evaluating the observational approach and suggesting additional tests to examine its robustness rather than coming up with alternative attribution approaches.
@@
The rest of the paper is structured as follows. Section 2 describes the statistical models, observation datasets, and methodologies. Section 3a shows that a GMST-dependence of distribution parameters does not necessarily represent the effects of anthropogenic climate change using an analysis of a CESM preindustrial simulation (i.e., without time-varying anthropogenic and natural forcings). We then further examine the three attribution cases in the remainder of section 3. Section 4 provides a summary and discussion of the results.
@@ 2.2 Generalized Pareto Distribution (GPD)
$GPD(x,\mu,\sigma,\xi)=\frac{1}{\sigma}\left(1+\xi\frac{x-\mu}{\sigma}\right)^{-(\frac{1}{\xi}+1)}.$ (6)
HW2022 assumed that the threshold parameter, which they denote $\mu$ (denoted $u$ by Coles et al., 2001) and scale parameter $\sigma$ both depend exponentially on the 4-year running-averaged GMST, $\overline{T}$ such that,
$\mu$ $=\mu_{0}\exp({\alpha\overline{T}}/{\mu_{0}}),$
$\sigma$ $=\sigma_{0}\exp({\alpha\overline{T}}/{\mu_{0}}).$ (7)
The exponential fit was assumed by HW2022 in order to ensure that the scaling factor $\exp(\alpha\overline{T}/\mu_{0})$ applied to both $\mu_{0}$ and $\sigma_{0}$ is strictly positive, so the distribution has fixed dispersion. We obtain the optimized GPD parameters ($\mu_{0},\sigma,\xi,\alpha$) by again maximizing the log-likelihood (Eq. 4.10, Coles et al., 2001),
$l(\mu,\sigma,\xi)=\sum_{t=1}^{m}\log(GPD(x_{t},\mu,\sigma,\xi)),$ (8)
using the trust-region constrained algorithm in the Scipy minimize Python package (‘trust-constr’ method).
The GPD requires the bound $\sigma>0$ and consistency relations of the form $x\geq\mu$ (for $\xi\geq 0$), and $\mu\leq x$ and $x\leq\mu-\sigma/\xi$ (for $\xi<0$). As the optimization iteratively attempts to search for the optimal parameter values, these constraints are occasionally violated, leading to complex numbers or NaNs (not a number) in the likelihood. The GPD optimization is notably more prone to such failures and sensitive to the initial guesses of $\alpha,\mu_{0},\xi,\sigma$ than the GEV optimization. In order to avoid such search failures, we replace complex and NaN terms in the log-likelihood sum with large penalty terms whenever they occur due to constraints being violated. In addition, if the optimization fails, we re-run the optimization again with new randomly selected initial guesses until convergence to a solution that maximizes log-likelihood and that satisfies the constraints occurs. We note that related solutions have been used to deal with GEV/GPD convergence issues by Robin and Ribes (2020).
@@ 2.4 Deviance Statistics (Likelihood Ratio Test)
In general, when fitting a model to data, if the number of fit parameters is increased, one expects a better fit because of the greater flexibility of the fitted model. Specifically, in the case considered here, adding a parameter $\alpha$ to include the GMST dependence of the location/threshold parameter will increase the maximum log-likelihood of the data. The salient question is whether the improvement in the fit exceeds that expected simply on account of the larger number of model parameters. To evaluate the improvement of the fit, Coles et al. (2001, theorem 2.7) recommend the use of a deviance statistic, also known as a likelihood ratio test. This test uses a measure, $D$, equal to twice the difference of the sum of log-likelihoods of the model with $k$ additional parameters minus that of the simpler model. Because more parameters lead to a better fit and thus higher log-likelihood, $D$ is non-negative. $D$ is then compared to a $\chi_{k}^{2}$ distribution to assess if the $k$ additional parameters significantly improve the likelihood. For example, with $k=1$, the 95th percentile of $\chi_{1}^{2}$ is equal to 3.8. If $D$ is larger than this threshold, we conclude that adding an additional parameter significantly improved the fit. It can then be inferred that the added parameter—in our specific case, the GMST dependence, $\alpha$—meaningfully explains features of the data. On the other hand, if the deviance statistic is not significant, one concludes that the addition of another parameter to the statistical fit is not justified by the data and, therefore, that this parameter cannot be used to draw conclusions about the data. Specifically, if one finds that the addition of $\alpha$ is not justified, the implication is that the estimated value should not be used to calculate the effects of climate change on the extreme events under examination. We also estimate the distribution and uncertainty of $D$ through bootstrap resampling of the data 5,000 times.

### v3

# On the attribution of weather events to climate change using empirically-fit extreme value distributions

@@ Abstract
Changes in extreme weather events are a potentially important aspect of anthropogenic climate change (ACC), yet, are difficult to attribute to ACC because the record length is often similar to, or shorter than, extreme-event return periods. This study is motivated by the “World Weather Attribution Project” (WWA) and, specifically, their approach of fitting extreme value distribution functions to local observations. The approach calculates the dependence of distribution parameters on global mean surface temperature (GMST) and uses this dependence to attribute extreme events to ACC. Applying this method to preindustrial climate simulations with no time-varying greenhouse gas forcing, we still find a strong dependence of distribution parameters on GMST. This dependence results from internal climate variability (e.g., ENSO) affecting both extreme events and GMST. Therefore, dependence on GMST does not necessarily imply an effect of ACC on extremes. We further consider for these three cases whether an extreme value, normal, or log-normal distribution better represents the data; if a GMST-dependence of distribution parameters is justified using a likelihood ratio test; and if a meaningful attribution is possible given uncertainties in GMST dependence. In one case we find that an attribution of Australia’s 2020–2021 Bushfires to ACC is difficult due to the effects of internal variability. In another involving the 2019–2021 drought in Madagascar we find that the small number of available data points precludes a meaningful attribution analysis. Overall, we find that the effects of internal climate variability on GMST and the uncertain relationship between GMST and regional extremes may lead to inaccurate attribution conclusions using the WWA approach.
@@ 1 Introduction
The Intergovernmental Panel on Climate Change Eyring et al. (2021); IPCC (2023, 2021) found that global mean surface temperature (GMST) between 2011–2020 was 1.09 C [0.95 C–1.20 C] above preindustrial (1850–1900) levels. Although climate change is clearly manifested in GMST warming, and there are reasons to expect extreme events to change with mean climate, it is a complex task to attribute individual extreme weather events to anthropogenic climate change (ACC). Extreme events—i.e., heatwaves, droughts, and heavy precipitation—are rare by definition, so the observational record may be too short to detect anthropogenic signals. This issue is especially pervasive in lower-income countries where data are sparse, and the historical record is short Otto et al. (2020). The situation is further problematic for extreme precipitation event attribution, which is highly sensitive to the specific timescales and percentiles used to define extreme precipitation events Pendergrass (2018).
@@
The field of attribution of extreme events has progressed significantly over the past decade and encompasses several approaches (see example reviews and cited references within, Stott et al., 2010; Hegerl and Zwiers, 2011; National Academies of Sciences, 2016; Clarke et al., 2022; Philip et al., 2020). The approaches that have been applied include the fraction of attributable risk Allen (2003); Stott et al. (2004), using simulations from climate models, using constraints derived from optimal fingerprinting techniques Christidis et al. (2015), using climate models to validate observed probability density functions (PDFs, Perkins et al., 2007; Lobeto et al., 2021), and the “Analogue-Based Approach” Lorenz (1969); Yiou et al. (2007); Vautard and Yiou (2009); Faranda et al. (2022). Another methodology fits statistical distributions to the historical record (again, either observed or from climate models) in what Stott et al. (2016) refer to as “Empirical-Based Approaches.” This approach assesses how the return period of particular events has changed over the historical record, with the purpose of detecting potential trends in the frequency of extreme events. These statistical approaches have been used to study trends in extreme events such as flooding in Thailand in 2011 Van Oldenborgh et al. (2012). Approaches using climate models allow for a better estimate of the magnitude and direction of human-induced contributions to an extreme event than is possible using observations alone.
@@
The climate model approaches detailed in the above paragraph can also be performed with Single Model Initial-condition Large Ensembles (SMILEs), composed of a climate model run several times with slightly different initial conditions. Averaging over multiple climate model ensemble members can reduce the signal from modes of internal variability, which are typically uncorrelated between ensemble members, facilitating the identification of forced climate change signals. This averaging generally improves the detection of trends in the magnitude or frequency of an extreme event, which are difficult to quantify in a single simulation or in the observational record where the signal from internal variability may be dominant Yoshida et al. (2017); Yamaguchi et al. (2020). These simulations have been used in the context of tropical cyclone slowdown Yamaguchi et al. (2020), as well as drought events in South and East Africa Lott et al. (2013); Pascale et al. (2020). The attribution literature is extensive, and for further detailed discussion, we refer the reader to the above reviews and to recent studies Stott et al. (2016); Knutson et al. (2017); Seneviratne et al. (2021).
@@
One particular approach that is the focus of this paper and that has been used in recent years by the WWA as a rapid response to extreme events attempts to fit the observed data to an extreme value PDF The World Weather Attribution project (2024). The fit allows for a dependence of the distribution parameters on explanatory variables, typically the GMST (e.g., Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022; Philip et al., 2020). A GMST dependence is then interpreted as a response of extreme events to ACC. This dependence is then used to attribute extreme events by quantifying the effects of anthropogenic forcing on the probability of their occurrence. Technically, this is achieved by writing the location parameter of the extreme event distribution, $\mu$, as a function of $\overline{T}$, the (low-pass filtered, via a 4-year running average) GMST “as a measure of anthropogenic climate change” Harrington et al. (2022), such that $\mu=\mu_{0}+\alpha\overline{T}$. The GMST-dependence, $\alpha$, is estimated via a fit to observations. The fitted PDF is then used to evaluate the probability of extreme events at 1900, $P(\mu_{0}+\alpha\overline{T}(1900))$ relative to those at 2020, $P(\mu_{0}+\alpha\overline{T}(2020))$. This allows for estimating how the probability of an extreme event changed between these times. A change in the probability of an extreme event is inferred by the WWA approach if $\alpha$ is non-zero and in proportion to the change in GMST.
@@
Our focus is on the empirically-based approaches used by the WWA, but it should be noted that WWA also requires evidence from climate model simulations before making any attribution statements. WWA often uses multiple climate models and is careful to exclude models based on their performance and return periods relative to observations (e.g., Pinto et al., 2023). In this study we use a single model (CESM) to demonstrate some of our points but focus on assessment of observational methodologies associated with fitting of extreme value distributions to observations. Our goal is limited to evaluating the observational approach described by WWA and suggesting additional tests to examine its robustness, as opposed to developing alternative attribution approaches.
Motivated by these pioneering efforts based on the fit to extreme event PDFs, we reexamine three test cases analyzed by WWA: the Siberian heat wave analysis of Ciavarella et al. (2021, hereafter CC2021), the analysis of high temperatures associated with the Australian Bushfire by van Oldenborgh et al. (2021, hereafter OK2021), and the Madagascar drought analysis of Harrington et al. (2022, hereafter HW2022). These cases were chosen as they represent a range of applications, including the attribution of heatwaves at high and low latitudes (Siberia vs. Australia) and of an extreme drought. Each of these studies leverages either a Generalized Extreme Value (GEV) or Generalized Pareto Distribution (GPD) fit to assess potential contributions to extreme events from ACC. WWA studies use both climate model simulations and observational analysis, and we focus on the observational component. In each of the three cases we use the CESM Large Ensemble Kay et al. (2015) to put the observed instrumental record into perspective, and we explore potential caveats concerning the methods examined and suggest how these limitations might be addressed.
The issue exposed by our analysis is simple: internal variability modes such as ENSO affect both local temperature extremes and GMST. In some cases, the effect on local extremes (say the effects of ENSO on heatwaves in Australia) may be large, but the effect on GMST is small. This might lead to a false conclusion that a small GMST signal leads to a comparatively large effect on local extremes. To demonstrate this point, we analyze a preindustrial climate model run with no anthropogenic forcing. We find, for example, that the fit of Australian heat extremes leads to values of $\alpha$ that are substantially greater than zero in the absence of ACC because of the effects of internal climate variability modes such as ENSO on both GMST and on local extremes. Our findings show that $\alpha$ can be strongly affected by internal variability, biasing the attribution results and complicating the WWA approach.
Our study addresses four questions: (1) Is a GMST dependence of the PDF parameters necessarily an indication of the signal of ACC? (2) Is a GEV or GPD necessary to fit the available observations, or are more standard normal or log-normal distributions sufficient? (3) Is the addition of GMST-dependent distribution parameters, which form the basis for the attribution analysis, justified statistically, and is the uncertainty range of these parameters, specifically $\alpha$, sufficiently small to allow for attribution? (4) What do SMILEs tell us about the amount of data needed to determine the GMST dependence with confidence? WWA uses multiple attribution tools, including analysis of simulations from multiple models and observations Philip et al. (2020). We focus only on the latter; specifically, on the fitting of extreme value distributions to observations. Our goals are limited to evaluating this observational approach and suggesting additional tests to examine its robustness, and do not include developing alternative attribution approaches.
@@
The rest of the paper is structured with Section 2 describing the statistical models, observation datasets, and methodologies. Section 3a shows that a GMST-dependence of distribution parameters does not necessarily represent the effects of ACC using an analysis of a CESM preindustrial simulation (i.e., without time-varying anthropogenic and natural forcings). We then further examine the three attribution cases in the remainder of section 3. Section 4 provides a summary and discussion of the results.
@@ 2.2 Generalized Pareto Distribution (GPD)
$GPD(x,u,\sigma,\xi)=\frac{1}{\sigma}\left(1+\xi\frac{x-u}{\sigma}\right)^{-(\frac{1}{\xi}+1)}.$ (6)
HW2022 assumed that the threshold parameter, which they denote $u$ (denoted $\mu$ by Philip et al., 2020) and scale parameter $\sigma$ both depend exponentially on the 4-year running-averaged GMST, $\overline{T}$ such that,
$u$ $=u_{0}\exp({\alpha\overline{T}}/{u_{0}}),$
$\sigma$ $=\sigma_{0}\exp({\alpha\overline{T}}/{u_{0}}).$ (7)
The exponential fit was assumed by HW2022 in order to ensure that the scaling factor $\exp(\alpha\overline{T}/u_{0})$ applied to both $u_{0}$ and $\sigma_{0}$ is strictly positive and maintain a fixed ratio. We obtain the optimized GPD parameters ($u_{0},\sigma,\xi,\alpha$) by again maximizing the log-likelihood (Eq. 4.10, Coles et al., 2001),
$l(u,\sigma,\xi)=\sum_{t=1}^{m}\log(GPD(x_{t},u,\sigma,\xi)),$ (8)
using the trust-region constrained algorithm in the Scipy Python package (‘trust-constr’ method).
The GPD requires the bound $\sigma>0$ and consistency relations of the form $x\geq u$ (for $\xi\geq 0$), and $u\leq x$ and $x\leq u-\sigma/\xi$ (for $\xi<0$). As the optimization iteratively searches for the optimal parameter values, these constraints are occasionally violated, leading to complex numbers or NaNs (not a number) in the likelihood. The GPD optimization is notably more prone to such failures and sensitive to the initial guesses of $\alpha,u_{0},\xi,\sigma$ than the GEV optimization. In order to avoid such search failures, we replace complex and NaN terms in the log-likelihood sum with large penalty terms whenever they occur due to constraints being violated. In addition, if the optimization fails, we re-run the optimization again with new randomly selected initial guesses until convergence to a solution that maximizes log-likelihood and satisfies the constraints is found. We note that related solutions have been used to deal with GEV/GPD convergence issues by Robin and Ribes (2020a).
@@ 2.4 Deviance Statistics (Likelihood Ratio Test)
In general, when fitting a model to data, if the number of fit parameters is increased, one expects a better fit because of the greater flexibility of the fitted model. Specifically, in the case considered here, adding a parameter $\alpha$ to include the GMST dependence of the location/threshold parameter will increase the maximum log-likelihood of the data. The salient question is whether the improvement in the fit exceeds that expected simply on account of the larger number of model parameters. To evaluate the improvement of the fit, Coles et al. (2001, theorem 2.7) recommend the use of a deviance statistic, also known as a likelihood ratio test. This test uses a measure, $D$, equal to twice the difference of the sum of log-likelihoods of the model with $k$ additional parameters minus that of the simpler model. Because more parameters lead to a better fit and thus higher log-likelihood, $D$ is non-negative. $D$ is then compared to a $\chi_{k}^{2}$ distribution to assess if the $k$ additional parameters significantly improve the likelihood. For example, with $k=1$, the 95th percentile of $\chi_{1}^{2}$ is equal to 3.8. If $D$ is larger than this threshold, we conclude that adding an additional parameter significantly improves the fit. It can then be inferred that the added parameter—in our specific case, the GMST dependence, $\alpha$—meaningfully explains features of the data. On the other hand, if the deviance statistic is insignificant, one concludes that the addition of another parameter to the statistical fit is not justified by the data and, therefore, that this parameter cannot be used to draw conclusions about the data. Specifically, if one finds that the addition of $\alpha$ is not justified, the implication is that the estimated value should not be used to calculate the effects of climate change on the extreme events under examination. We also estimate the distribution and uncertainty of $D$ through bootstrap resampling of the data 5,000 times.